OpenStudy (anonymous):

I need to write this expression into a single logarithmic form: \[2 log_b q + 8 log_b t\] I used the property of that: \[log_m^n = n log m\] So would the answer be: \[(2 + 8) log_b (q + t)\]

hartnn (hartnn):

nopes, lets revise the log properties

hartnn (hartnn):

\(\large A\log B = \log B^A \\\ \large 2 \log q = .... ? \)

hartnn (hartnn):

and 8 log t = ... ?

OpenStudy (anonymous):

\[\large 2 \log q = \log_q 2\] \[\large\log_t 8\]

hartnn (hartnn):

oh, do you mean

hartnn (hartnn):

 ?